# Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits

### v1

@@
Abstract
This paper is motivated by recent developments in the linear bandit literature, which have revealed a discrepancy between the promising empirical performance of algorithms such as Thompson sampling and Greedy, when compared to their pessimistic theoretical regret bounds. The challenge arises from the fact that while these algorithms may perform poorly in certain problem instances, they generally excel in typical instances. To address this, we propose a new data-driven technique that tracks the geometry of the uncertainty ellipsoid, enabling us to establish an instance-dependent frequentist regret bound for a broad class of algorithms, including Greedy, OFUL, and Thompson sampling. This result empowers us to identify and “course-correct” instances in which the base algorithms perform poorly. The course-corrected algorithms achieve the minimax optimal regret of order $\widetilde{\mathcal{O}}(d\sqrt{T})$, while retaining most of the desirable properties of the base algorithms. We present simulation results to validate our findings and compare the performance of our algorithms with the baselines.
@@ 1 Introduction
Multi-armed bandits (MABs) have garnered significant attention as they provide a well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems. In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards. This paper specifically focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases. In this model, the system parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield rewards equal to the inner product of $\theta^{\star}$ and the chosen action, along with mean-zero, subgaussian noise. The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time. The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total rewards achieved by the policy and the maximum achievable reward. Achieving this objective necessitates striking a balance between exploration and exploitation. In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards. Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985). Other popular strategies include Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions. Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
@@
In the linear bandit setting, two types of regret are considered. The Bayesian regret is applicable when the parameter $\theta^{\star}$ is random, and the regret is averaged over three sources of randomness: the observation noise, the randomized algorithm, and the random parameter $\theta^{\star}$. Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al. (2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to logarithmic factors. On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of noise and algorithm.
@@
In the frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al. (2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound. However, for a frequentist variant of LinTS referred to as TS-Freq, Agrawal and Goyal (2013) and Abeille et al. (2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret by inflating the posterior distribution. This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$. Recent work by Hamidi and Bayati (2020a) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved. While no general theoretical guarantee exists for the Greedy algorithm in linear bandit problems, it is known that both Greedy and LinTS perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
@@
Despite the theoretical gaps in performance, LinTS has shown strong empirical performance, indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently. Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021). Additionally, while OFU-based algorithms are computationally expensive (generally NP-hard as discussed by Dani et al. (2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency. This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify problematic instances where LinTS and Greedy could potentially fail, and apply a “course-correction” to ensure competitive frequentist regret bounds? And can this be achieved without compromising their impressive empirical performance and computational efficiency? In this paper, we provide positive answers to both questions. Specifically, we make the following contributions.
  • Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid. This technique is essential in maximizing the utilization of historical data, thereby enabling the “course-correction”.
  • We derive a general instance-dependent frequentist regret bound for the proposed POFUL algorithms, a family of algorithms including OFUL, LinTS, TS-Freq, and Greedy as special cases. The bound can be efficiently calculated in a data-driven manner based on the observation history.
  • We propose variants of LinTS and Greedy that enjoy the minimax optimal frequentist regret, while retaining most of the desirable properties such as computational efficiency.
@@
The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid. This is in stark contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their overly narrow concentration on comparatively rudimentary summary statistics.
Related literature.
The ideas presented in this paper bear similarities to existing literature that explores different forms of exploration (Russo and Van Roy, 2016; Kirschner and Krause, 2018) or recent works by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b) that aim to utilize observed data for reducing exploration. However, the methodology and the manner in which the data is utilized in this paper differ significantly. For instance, Bastani et al. (2021) examines the minimum eigenvalue of the sample covariance matrix which is a single-parameter summary statistic of the observed data, while Hamidi and Bayati (2020b) incorporates information from the (one-dimensional) confidence intervals around the reward values. In contrast, our approach leverages the complete geometric information of the $d$-dimensional confidence ellipsoid.
@@ 2.1 Assumptions
The action sets $\{\mathcal{X}_{t}\}$ are uniformly bounded and closed subsets of $\mathbb{R}^{d}$, such that $\|x\|\leq X_{t}$ for all $x\in\mathcal{X}_{t}$ and all $t\in[T]$, where $X_{t}$’s are known and $\sup_{t\geq 1}\left\{X_{t}\right\}<\infty$. Without loss of generality, we normalize the actions such that $\mathcal{X}_{t}\subset{\mathcal{S}}_{d-1}$ for all $t\in[T]$.
@@ 2.2 Regularized Least Square and Confidence Ellipsoid
Let $(x_{1},\dots,x_{t})$ be an arbitrary sequence of actions and $(r_{1},\dots,r_{t})$ be the corresponding rewards. In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator. Fix a regularization parameter $\lambda$, the design matrix, and the Regularized Least Square (RLS) estimate are respectively given by
@@
Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that
$\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta),\quad\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$
for all $x\in\mathbb{R}^{d}$, where
@@
The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret. Basically, this proposition is saying that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
@@ 3 POFUL Algorithms
In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$. The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability. This turns into the following condition for $\{\iota_{t}\}_{t\in[T]}$ and $\{\widetilde{\theta}_{t}\}_{t\in[T]}$.
@@
The intuition behind POFUL is combining the exploration mechanism of OFUL and LinTS. OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores by sampling from the confidence ellipsoid. We allow POFUL to select an arbitrary pivot from $\mathcal{E}_{t}^{PVT}$ to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$. By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
@@
LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$. We remark that setting the inflation parameter $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm. On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al. (2017), denoted by TS-Freq. This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
@@ 4 Frequentist Regret Analysis of POFUL
$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator. Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate. By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda)$. Hence the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which are characterized by the confidence ellipsoids $\mathcal{E}^{RLS}_{t}$. $\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$. Similar to $\mathcal{R}^{RLS}(T)$, the upperbound is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$. As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$. Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound. For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$. However, later we see this could be problematic when bounding $\mathcal{R}^{PE}(T)$.
@@
$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term. In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$. In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$. For OFUL, this term is bounded properly with high probability by the construction of OFUL actions. For LinTS, the only known analysis due to Agrawal and Goyal (2013); Abeille et al. (2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$. The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$. In Abeille et al. (2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e. $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$. For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures. The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy.

### v3

@@
[…]
@@ 1 Introduction
Multi-armed bandits (MABs) have garnered significant attention as they provide well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems. In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards.
This paper focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases. In this model, the problem parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield noisy rewards with mean equal to the inner product of $\theta^{\star}$ and the chosen action. The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time. The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total expected rewards achieved by the policy and the maximum achievable expected reward.
@@
Achieving this objective necessitates striking a balance between exploration and exploitation. In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards. Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985; Auer, 2002; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). Another popular strategy is Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions. Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
@@
In the linear bandit setting, two types of regret are considered. The Bayesian regret is applicable when the parameter $\theta^{\star}$ is random, and the regret is averaged over three sources of randomness: the observation noise, the randomized algorithm, and the random parameter $\theta^{\star}$. Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al. (2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to polylogarithmic factors.
@@
On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of the noise and the algorithm. In this frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al. (2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound. However, for a frequentist variant of LinTS referred to as TS-Freq, which inflates the posterior distribution, Agrawal and Goyal (2013) and Abeille et al. (2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret. This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$. Recent work by Hamidi and Bayati (2020b) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved. For the Greedy algorithm in linear bandit problems, no general theoretical guarantee exists. Moreover, it is known that both LinTS and Greedy might perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
@@
Despite the theoretical gaps in performance, LinTS has shown strong empirical performance (Russo et al., 2018), indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently. Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021). Additionally, while OFU-based algorithms are computationally expensive (generally NP-hard as discussed by Dani et al. (2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency. This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify, in a data-driven fashion, problematic instances where LinTS and Greedy could potentially fail and apply a “course-correction” to ensure competitive frequentist regret bounds? And can this be achieved without compromising their impressive empirical performance and computational efficiency? In this paper, we provide positive answers to both questions. Specifically, we make the following contributions.
  • We develop a real-time geometric analysis technique for the $d$-dimensional confidence ellipsoid surrounding $\theta^{\star}$. This method is crucial for maximizing the use of historical data, advancing beyond methods that capture only limited information from the confidence ellipsoid, such as a single numerical value. Consequently, this facilitates a more precise “course-correction”.
  • We introduce a comprehensive family of algorithms, termed POFUL (encompassing OFUL, LinTS, TS-Freq, and Greedy as specific instances), and derive a general, instance-dependent frequentist regret bound for them. This bound is efficiently computable using data observed from previous decision epochs.
  • We introduce course-corrected variants of LinTS and Greedy that achieve minimax optimal frequentist regret. These adaptations maintain most of the desirable characteristics of the original algorithms.
@@
The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid. This is in contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their focus on summary statistics.
1.1 Other Related literature
We note that Valko et al. (2014); Kocák et al. (2014, 2020); Kocák and Garivier (2020) consider the spectral of graphs and study Spectral Bandits. In their work, the rewards of arms are represented as smooth functions on a graph, allowing the exploitation of a low-rank structure for improved algorithmic performance and a regret guarantee that is independent of the original number of actions. In contrast, our approach leverages the spectral properties of the sample covariance of actions, which is distinct from graph spectral analysis. Additionally, our research addresses the broader context of stochastic linear bandit problems without presupposing any low-rank structure.
The concepts in this paper bear similarities to existing research that investigates various exploration strategies (Russo and Van Roy, 2016; Kirschner and Krause, 2018), as well as recent studies by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b, a) that focus on using observed data for reducing exploration. Nonetheless, our methodology and the specific use of data markedly differ. For example, Bastani et al. (2021) emphasizes the minimum eigenvalue of the sample covariance matrix, essentially a single-parameter summary of the observed data. On the other hand, Hamidi and Bayati (2020a) uses information from the one-dimensional confidence intervals surrounding the reward values. A more closely related paper, Hamidi and Bayati (2020b), uses spectral information to enhance the performance of TS in LB problems. They look at a single summary statistics of the confidence set that they call thinness coefficient, to decide if the posterior of TS should be inflated. In contrast, with these studies, our approach capitalizes on the full geometric details of the $d$-dimensional confidence ellipsoid, utilizing richer geometric information.
1.2 Organization
The structure of this paper is as follows: Section 2 introduces the notations and problem formulation. In Section 3-4, we present the POFUL family of algorithms and analyze their frequentist regret. Sections 5-6 detail our data-driven approach and meta-algorithm for course-correction. Section 7 discusses simulations using synthetic and real-world datasets. The paper concludes in Section 8 with a summary of findings. All technical proofs, further discussions, and supplementary materials are provided in the appendices.
endcomment
@@ 2.1 Assumptions
The action sets $\{\mathcal{X}_{t}\}$ are uniformly bounded and closed subsets of $\mathbb{R}^{d}$, such that $\|x\|\leq X_{t}$ for all $x\in\mathcal{X}_{t}$ and all $t\in[T]$, where $X_{t}$’s are known and $\sup_{t\geq 1}\left\{X_{t}\right\}<\infty$.
@@ 2.2 Regularized Least Square and Confidence Ellipsoid
Let $(x_{1},\dots,x_{t})$ be an arbitrary sequence of actions and $(r_{1},\dots,r_{t})$ be the corresponding rewards. In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator. Fix a regularization parameter $\lambda$, the sample covariance matrix, and the Regularized Least Square (RLS) estimate are respectively given by
@@
Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that $\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)$ and $\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$ for all $x\in\mathbb{R}^{d}$, where
@@
The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret. Essentially, this proposition suggests that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
@@ 3 POFUL Algorithms
In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). It allows us to analyze the frequentist regret of common algorithms in a unified way.
At a high level, POFUL is designed to encompass the exploration mechanism of OFUL and LinTS. POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$. The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability. This turns into the following condition for $\{\iota_{t}\}_{t\in[T]}$ and $\{\widetilde{\theta}_{t}\}_{t\in[T]}$.
@@
Recall OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores through random sampling within the confidence ellipsoid. We let POFUL select an arbitrary pivot (which can be random) from $\mathcal{E}_{t}^{PVT}$ and maximize the optimistic reward to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$. By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
@@
LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$. We remark that setting the inflation parameter $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm. On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al. (2017), namely TS-Freq. This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
@@ 4 Frequentist Regret Analysis of POFUL
$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator. Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate. By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda)$. Hence the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which are characterized by the confidence ellipsoids $\mathcal{E}^{RLS}_{t}$. $\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$. Similar to $\mathcal{R}^{RLS}(T)$, the upperbound is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$. As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$. Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound. For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$. However, later we see this could be problematic for bounding $\mathcal{R}^{PE}(T)$.
@@
$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term. In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$. In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$. For OFUL, this term is bounded properly with high probability by the construction of OFUL actions. For LinTS, the only known analysis due to Agrawal and Goyal (2013); Abeille et al. (2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$. The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$. In Abeille et al. (2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e. $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$ with a constant probability. For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures. The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy that cannot be analyzed by the optimism-based methods.